OpenStudy (anonymous):

What is the effective annual interest rate of an investment that pays 5.2% annual interest, compounded quarterly?

OpenStudy (dumbcow):

\[\rightarrow (1+\frac{.052}{4})^{4} -1\] in general \[r = (1+\frac{i}{n})^{n} -1\]

OpenStudy (anonymous):

Oh wow... ok so where i get confused is plugging in the numbers correctly

OpenStudy (anonymous):

hold on a sec let me try to work it out and see here

OpenStudy (dumbcow):

just use the general formula as a reference

OpenStudy (anonymous):

so first I will divide .052/4 right? then by 4 correct??? this is where I get totally lost...thanks by the way for helping me Im really having a hard time with these problems

OpenStudy (dumbcow):

ok divide .052 by 4 add 1 raise to 4th power subtract 1 :)

OpenStudy (anonymous):

so its (1.013)(4) -1???

OpenStudy (anonymous):

3.052?

OpenStudy (anonymous):

that cannot be right lol

OpenStudy (dumbcow):

no 4th power not multiply by 4

OpenStudy (anonymous):

oh crap hang on lol

OpenStudy (anonymous):

UGHHHHH what am I doing wrong I keep getting the same stupid answer I know not to the 4th power got that but not getting any of the answers supplied ughhh

OpenStudy (anonymous):

(1.013) ok so then its 1.013 4 times?? why am I confusing myself I have been doing this crap since 3 this afternoon lol

OpenStudy (anonymous):

not multiply though

OpenStudy (anonymous):

OH no its 5.30% is that right??

OpenStudy (dumbcow):

(1.013)^4 = 1.053 yes you got it now :)

OpenStudy (anonymous):

woohoo awesome!!! Can you help me with anymore or do you have time?

OpenStudy (dumbcow):

sure

OpenStudy (anonymous):

OH THANK YOU!!!!! ;) Ok so this is another..... $4200 is invested for 3 months at an annual simple interest rate of 13%. How much interest will be earned?

OpenStudy (anonymous):

so I got P= 4200 R= .013

OpenStudy (anonymous):

T=3?

OpenStudy (dumbcow):

oh simple interest is easy \[I = P*r*t\]

OpenStudy (dumbcow):

"t" is always in years...we want 3 months what fraction of a year is that ?

OpenStudy (anonymous):

3/365??? that maybe wrong?

OpenStudy (dumbcow):

haha no 3/12 = 1/4

OpenStudy (dumbcow):

12 months not 365 days

OpenStudy (anonymous):

LOL I feel so dumb haha!!!! I was thinking 365 days in a year lol

OpenStudy (anonymous):

so then do I change 1/4 into decimal?

OpenStudy (dumbcow):

if you want doesnt matter

OpenStudy (anonymous):

oh ok just sec then

OpenStudy (anonymous):

So I get 13.65 but thats not one of my answers

OpenStudy (anonymous):

what did I do wrong should be right?

OpenStudy (anonymous):

OH no 136.50 is it

OpenStudy (dumbcow):

i think it should be 136.50 rate = 13% = 0.13

OpenStudy (dumbcow):

your decimal point was off :)

OpenStudy (anonymous):

woohooo!!! yeah whoops lol

OpenStudy (anonymous):

Have a couple more you have time dont want to bug you

OpenStudy (dumbcow):

i got 10 min

OpenStudy (anonymous):

Calculate the compound amount using the compound amount formula when P= 1,600, r=8% compounded semiannually, and t=14 years. Round to the nearest cent. OK I WILL try to hurry

OpenStudy (anonymous):

P=1600 R=.08 T=14 so is it just multiplying?

OpenStudy (anonymous):

OH wait it says semiannually ok so thats not just all hu

OpenStudy (anonymous):

You there? :)

OpenStudy (dumbcow):

semi-annual means compounded 2 times a year...n=2 use the general formula \[P(1+\frac{r}{n})^{nt}\]

OpenStudy (anonymous):

OH CRAP

OpenStudy (anonymous):

whats n again?

OpenStudy (anonymous):

2? semiannualy right??? no I dont know????

OpenStudy (anonymous):

oh I see where you put that hold on

OpenStudy (anonymous):

so far I have 1600(1.04)28th power right??

OpenStudy (dumbcow):

correct, just plug that into calculator gotta go

OpenStudy (anonymous):

crap not getting it 1.04 28th i dont times it right what do i do?

OpenStudy (dumbcow):

1.04^28

OpenStudy (dumbcow):

its an exponent

OpenStudy (anonymous):

k thanks for all your help alot!!!!

OpenStudy (dumbcow):

yw, have a good night

OpenStudy (anonymous):

u still there by chance sorry...is it 4,699.51 just to make sure :)